Find the average rate of change of f(x)=3x^2-2, from 2 to 4.
Answer #1

f(x)=309_2 ato¢ The average rate of change f(a) f(a) - (344)2-a) - (36a)&na) 4-8 (3x16-2) = (3x4-2) = (48-a) -(18-a) 46 - 10